Systems Staff Meeting Notes 9/21/98

Systems Staff Meeting Notes
September 21, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Suzy Shaw

Announcements
  1. The documentation for the IR Database was found in Winston's office. If anyone finds anymore, please give it to Bobbie Meng.
  2. ITSA Day (Information Technology Security Awareness) - Thursday, October 1, 8AM-5PM at the Reitz Union. Dale would like the liaisons to attend, Bill will be going & he encourages anyone interested to do the same. Security policies will be discussed.
  3. Outcome of the lawsuit against the hacker in the CIS ceiling this past summer - Lombardi is now interested in whom has access to the ufl.edu domain. He wants a list of names & access capabilities. Plans are to run security screens on all servers on campus - to hammer on them to see if they can break in. Campus has been scanned from the outside, the administration is looking for volunteers to have their servers scanned. This will mean that we'll have to tighten up on our security, permission & passwords.
  4. Positions: A&P Coordinator - Barbara Oliver & Winston rescheduled their meeting to go over the resumes to Wednesday. There has been some confusion as to who's in the pool - Bill would like to see the list.
  5. Sharon Henry's old position will be resubmitted as is.
  6. Lee's old position - will try to re-title to Computer Repair Technician. The rate is available, unless it gets knocked down. Will post it for two weeks, it will be a minimum of five weeks until it is filled.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Email server - Compaq monitor received. Compaq can't build the CPU, they have no memory chips or disk drive. That project is on hold for the moment.
  2. Modem for Marcia Pearce's laptop was just received this morning. The mailroom received it Fed Ex - they don't call us immediately to let us know - they wait to receive the PO. Bobbie just happened to see it when she went over there to pick up the floppy cables that had come in.
  3. Floppy cables received.
  4. Bobbie points out that items bought on a blanket account do not generate any paperwork indicating their destination. Bobbie wanted to set up a file on NT2, Bill says Scott Olipra should generate letters to that effect.
  5. We should be receiving 2 ethernet adapters for laptops. One is for Rich, one for the shelf.
Networking
  1. Winston noted that sign-ons to email were disappearing this morning in Preservation, Cataloging & Resource Services. The switch they all share appeared to be normal - NERDC could ping it, but it was not forwarding traffic to the outside world. Bill did a soft reboot - then had to power cycle - came back up in about 2 minutes. The switch is not talking to the management console - it may be failing. We do have a spare switch intended for the classroom. It has two copper feeds, would have to use the media converter.
  2. These problems are probably totally independent of the other difficulties that have been occurring in Preservation. They had a number of misconfigured work stations, duplicate IP's and interrupt conflicts.
  3. Room 100 Walk-Up Access - probably won't open until next week as the Capitol Development Campaign will override the use of the room.
Software Projects Status
  1. Winston worked on the glitch in the Trouble Ticketing System - not as slow now, updated version.
  2. LTMS - Winston has been working on the LTMS viewer access for Reference. Bill says this is not necessary, they should be able to look up machine readable files on NOTIS. Need to make sure the TLA numbers are being entered into NOTIS.
  3. Personnel has been using the permanent employee side of the database. Winston told them they could keep the Mac until the end of the month. At that point, both Sharon's & Personnel's Macs will be ready for PKY.
  4. Brian has been working on LASSIE, should be done before October.
  5. Brian has been optimizing features to make the Special Collections database run better. So far, no problems - has been installed on two machines & is multi-user.
  6. Winston found a way to manually override DELPHI problem of opening & closing sessions in the wrong order.
  7. Timer software for John Ashcraft - no news.
  8. Barbara Oliver called Winston about the CLIO system used by David Fuller on a one month trial basis to replace ILL - the month is up & he wants to keep it - cost is $2000. Bill says it couldn't completely replace the system we designed for ILL. For OCLC transactions, it allows them to make one entry, but they still need the ILL tracking database written by Systems - it will save them time not rekeying.
DP Coordinator
  1. The yearly Compustat tapes have been received (20). This will be the first time to log them in & copy them under the new system. Debra will probably be working with Winston on this.
  2. Debra received an OCLC tape - unusual, not one she had received before. Back in the 80's there was discussion of using OCLC tapes for RLG loads. It is probably a database record update. It is a small 6" reel - probably not containing too much data. Rich will figure out if we have a use for it, may end up getting rid of it.
  3. Barbara Oliver called Debra asking for a trackball for Glenda - in the past, the policy was to have an ergonomic evaluation done first. Bill has a number of trouble tickets pending on this particular topic. Bill will try to get a clarification on the policy.
Calendar
  1. Bill out Sept. 23.
  2. Rich out Sept. 23-24 - Government Health Conference in Tallahassee.
  3. Brian out Oct. 2 & 5.
  4. Brian out Nov. 20-23.


Back to the Systems Homepage

Last updated February 24, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu